Suppose that labor demand is given by w = 80 - L, while labor supply is given by w = L. Suppose that, during a recession, labor demand decreases to w = 50 - L. If there are sticky wages, how many people become unemployed?


1
Expert's answer
2021-04-18T19:25:50-0400

Solution:

First find labor market equilibrium:

At equilibrium: Labor demand = Labor Supply

Labor demand = 80 – L

Labor supply = L

Therefore:

80 – L = L

80 = L + L

80 = 2L

L = 40

At equilibrium, labor units are 40.

Supply will be: 40

When demand reduces to 50 – L, the new equilibrium will be as follows:

Therefore:

50 – L = L

50 = L + L

50 = 2L

L = 25

The new equilibrium due to a fall in labor demand will be 25 units

Supply = 25


At this situation, we expect companies to reduce wages rather than sack employees during a recession.

Sticky wages mean that workers earnings do not change or adjust quickly to changes in the labor market conditions due to various reasons. Therefore, during a recession, employees wages will not be able to be reduced due to sticky wages. As such, the company will fire a few employees until the labor market equilibrium is attained.

The number of people who will become unemployed will be 15 employees.

That is: Old supply - new supply

40 – 25 = 15